# The Authentication Gap: Higher Education’s Widespread Noncompliance with NIST Digital Identity Guidelines

### v1

@@ 4.2 Data Collection
The data collection for this study followed a straightforward protocol for all selected institutions. Data collection took place between September 2023 and August 2024.
4.2.1 Data Format
We first identified the specific digital authentication standards from NIST SP 800-63-3 listed in Section 3 as the focus of our study. We then created a spreadsheet to support standardized data collection across standards and institutions. For each standard, we sought to answer the following question for each institution and code the answer as one of a limited set of data codes:
  1. 1.
    Does the practice associated with this standard appear in the publicly available online policies of the institution?
    1. (a)
      “No Policy Found": There is no information about this specific practice
    2. (b)
      “In Policy”: There is information about this specific practice
If the answer to question 1 was “in policy,” we then sought to answer the following question:
  1. 2.
    What is the institutional policy regarding this practice?
    1. (a)
      “Required”: The practice is required for all affiliates of the institution
    2. (b)
      “Required for Specific Affiliates”: Some affiliates of the institution are required to follow the practice. In this case, we also recorded which specific affiliates are subject to this requirement.
    3. (c)
    4. (d)
    5. (e)
We chose the answer codes for both questions after a trial data collection with three institutions. The codes covered all cases we found throughout the collection process. In addition to this consistently formatted data, we also took open-ended notes on any particularly interesting policies we discovered.
4.2.2 Data Sources
We sought to answer questions 1 and 2 above based on any publicly available online policies posted on official institutional websites. We used several methods to locate these policies, including (but not limited to):
@@ 4.2.2 Data Sources
This process was intentionally open-ended, reflecting the variety of institutional website structures and information posting practices. We found relevant policies most frequently in the following documents:
@@
    An “information security policy,” “cybersecurity policy,” or “IT policy” often posted by the IT department of the institution
@@
    IT “support pages” (also referred to as “guides,” “help desk,” “knowledge base,” or similar) posted by the IT department of the institution
  3. 3.
    A “privacy policy” or “privacy notice” often posted on the homepage of the institution
At least one author and a trained research assistant manually reviewed each of the relevant policies, recording answer codes for each standard.
4.3 Data Confirmation
Publicly available online information posted by institutions is not always up to date with actual practices. We therefore reached out to representatives at all studied institutions according to the below protocol for confirmation and/or updates to the information we gathered.
We first identified a contact email address for the information technology department at the institution. For most institutions, this was the email to an IT help desk or equivalent (e.g., “ithelp@” or “servicedesk@”). Some institutions had specific email addresses for security-related contacts, which we used if available (e.g. “information.security@”). Finally, a few institutional web sites directed queries directly to a specific IT staff member.
We then emailed all institutions with a customized email including an introduction to our study, a summary of the information we had already collected about the institution from their publicly available policies, and a request to complete a Google Form with any corrections, clarifications, or completions of missing data. We indicated that such responses would be appreciated and would ensure that the institution was not misrepresented in our paper.
Out of the 101 institutions we included in this study, only 13 responded to our outreach. Of these, 9 said that they could not provide any information for our study. Typical responses were akin to “Unfortunately we are not able to participate in this project” and “ITS will not respond to requests outside of their institution.” Only 4 institutions actually provided useful information by completing the Google form.
Most of these form responses corroborated the data we had already collected from the institutions’ online policies. In the case of discrepancies, we report the data from the online policies (rather than the form responses). This paper therefore provides a snapshot of publicly available online policies across the sector, unbiased by form responses from a small number of institutional representatives (who are incentivized to provide socially correct descriptions of their security practices).
@@ 4.4 Research Ethics
The Colgate University Institutional Review Board determined that this study was exempt. We only report data from publicly available sources posted online by the institutions we study. Although these data are not private, we anonymize our results, refraining from naming-and-shaming any particular institutions.
As previously discussed, the limited amount of data provided directly by the institutions for data confirmation purposes was not included in our analysis. We do not collect or report any information about individual affiliates at any institution.
@@ 4.5 Limitations
First, it is essential to remember that this study examines institutional policies, and that these policies may or may not represent actual, technologically-enforced institutional practices. For example, a policy might say that multi-factor authentication is required for all institution affiliates, but some affiliates might, in practice, be able to disable MFA via their internal profile settings. Similarly, a policy might state that yearly password expiration is required for all affiliates, but the current IT leadership might not actually enforce regular password cycling. Nevertheless, we expect posted policies to generally align with actual practices, especially since posted policies are often primarily intended as guides for affiliates (e.g., on IT “help” websites). The policies signal the intentions and values of the institution and influence the actions of affiliates and peer institutions. On a practical level, determining actual technically-enforced practices would require institutional affiliation or direct input from institutional representatives, which would be challenging given the low response rate to our data confirmation requests and hostility to discussing cybersecurity practices in most of the responses we did receive.
Second, the 101 institutions we studied represent only a fraction of the higher education sector. As noted above, we encourage follow-up studies examining digital authentication practices and other forms of cybersecurity compliance across other countries and institutional categories. The institutions in this paper are at the top end of prestige in their respective categories, which must be remembered when interpreting the results. These institutions, in contrast to less prestigious peers, are more likely to be well-resourced, making them more likely to have sufficient staff with expertise and availability to ensure the institution’s digital authentication (and other cybersecurity) practices follow accepted guidelines. The fact that we find many cases of noncompliance with NIST standards among these prestigious institutions’ online policies suggests that the overall rate of noncompliance is substantially higher across the higher education space.
Third, this study reflects a snapshot in time, with data collected between September 2023 and August 2024. Follow-up longitudinal studies are necessary to see whether the trajectory of institutional policies and practices is trending toward compliance with NIST standards.
@@ 5 Results
This section presents the results from our investigation of the publicly available online policies of 101 institutions of higher learning in the United States and Canada.
For each authentication practice described in Section 3, we first report what fraction of institutions in each category refer to the authentication practice in their online policies. We then focus on the subset of institutions with such information and report the prevalence of specific practices. We note which practices are in or out of compliance with NIST SP 800-63-3 and provide anecdotal examples.
Figure 1: Online publicly available policies of 101 higher education institutions regarding standards from NIST SP 800-63-3. Policies in green align with the standards. Policies in red are noncompliant.
@@ 5.1 Password Expiration
We found that requiring or recommending password expiration/cycling is still widespread across higher education (Figure 1). 47% (47) of all investigated institutions referred to password expiration in their online policies. Of these, 47% (22) required regular password expiration for all affiliates, 17% (8) required regular password expiration for specific affiliates (faculty/staff, administration, or “HIPAA affectees”), and 28% (13) recommend (but did not require) regular password expiration for all affiliates. These policies are in direct violation of the NIST 800-63-3 standard that password expiration should only occur in response to a known breach. Only 11% (5) of institutions that referred to password expiration in their online policies did so to explicitly disallow or discourage the practice.
The institutions that require or recommend password expiration for some or all affiliates do so for a wide range of expiration frequencies (Table 2). One year was the most common password expiration frequency we observed, but specific frequencies ranged from expiration every 1 month to every 400 days, with one institution recommending password expiration “often,” one recommending “periodically,” and several not listing a frequency in their publicly available policy.
@@
1, 3, or 6 months 1
@@
90 days 3
3 months 2
3-6 months 1
120 days 4
@@
Periodically 1
Often 1
Table 2: Password expiration frequencies for institutions that require or recommend regular password cycling.
Some of these institutions purposefully disregard the NIST standard. One university has a Q&A on “Why Are We Implementing A Password Change Initiative When It Is No Longer Considered Best-Practice?” that states
@@
A smaller percentage of US regional colleges (37%) and Canadian universities (40%) referred to password expiration in their policies than institutions in other categories ($\geq$60%). However, across institutions with policies referring to password expiration, a greater fraction of US regional colleges required expiration for all affiliates (67%) versus institutions in other categories ($\leq$50%).
@@ 5.2 Knowledge-based Authentication
The use of security questions (e.g. “What is the name of your first pet?”) during the authentication process remains alarmingly widespread across higher education (Figure 1). 45% (45) of all investigated institutions referred to the use of security questions during account creation, login, and/or recovery in their online policies. Of these, 62% (28) required security questions for all affiliates, 25% (11) required security questions for specific affiliates, and 11% (5) recommended (but did not require) security questions for all affiliates. All such policies are in violation of the NIST SP 800-63-3 standard that security questions are no longer recommended for any digital authentication process due to their inherent insecurity. Only a single institution discouraged the use of security questions.
Overall, a smaller percentage of US regional colleges (30%) referred to security questions in their policies than institutions in other categories (>>>52%). Across institutions with policies referring to security questions, a smaller percentage of US universities actually require them for all affiliates (33%) as compared to institutions in other categories (>>>69%).
@@
The institutions that use security questions typically do not provide any guidance about which questions or what types of answers are likely to be more secure. One institution described transitioning from requiring one security question to requiring three security questions. While this approach may be slightly more secure, it is still out of compliance with the NIST standard.
5.3 Password Composition Rules
Despite the NIST SP 600-83-3 standard that institutions should not impose password composition rules because they “do not significantly improve the security of selected passwords” [32], the presentation and enforcement of password composition rules appears widespread across higher education (Figure 1). 73% (74) of all investigated institutions referred to password composition rules in their online policies. Of these, 77% (57) require that affiliates meet minimum password composition rules during the password creation process. Another 22% (17) recommend (but do not require) that affiliates meet password composition rules.
Similar to previous cases, a smaller percentage of US regional colleges (58%) referred to password composition rules than institutions in other categories ($\geq$80%). Across institutions with policies referring to password composition rules, greater than 80% of US institutions across categories enforce compliance with these rules during password creation. A much smaller percentage of Canadian universities referring to password composition rules explicitly require compliance (41%). Figure 2 shows a representative example of password composition rules from a US regional college.
5.4 Password Strength Meters
Rather than specific password composition rules, NIST SP 800-63-3 recommends that institutions provide graphical password “strength meters” during the password creation process. Unfortunately, only 6% (6) of all investigated higher education institutions referred to password strength meters in their online policies (Figure 1). This low fraction is effectively uniform across institution categories. Of this fraction, one US national university and one US regional college require that affiliates meet some minimum strength requirement on a meter during password creation (e.g., a rating of “strong” on the meter shown in Figure 3 from a US regional college).
Figure 3: Example strength meter from a US regional college that requires a rating of “strong” during password creation.
5.5 Multi-factor Authentication
The use of multi-factor authentication (MFA) is widespread in higher education (Figure 1). 76% (77) of all investigated institutions referred to MFA in their online policies. Of these, 73% (56) required MFA for all affiliates, in compliance with NIST SP 800-63-3. A further 9% (7) required MFA for specific affiliates, and 16% (12) recommend (but did not require) MFA for all affiliates. The institutions that only require MFA for certain affiliates do so for employees only (i.e., not students) or only for IT staff.
A considerably smaller percentage of US regional colleges referred to MFA in their online policies than institutions in other categories (51% vs >>>90%). Of the institutions with policies referring to MFA, US national universities have the highest requirement rate for all affiliates (84%), followed by US liberal arts colleges (76%), Canadian global universities (71%), and US regional colleges (64%).
NIST SP 800-63-3 further specifies that email should not be considered as a valid second factor for multi-factor authentication. 91% (92) of the institutions did not provide any information about the use of email as a second factor in their online policies. Fortunately, many of these institutions use a third-party service for MFA, such as Duo, that does not allow email as a second factor [15]. However, 7% (7) of investigated institutions do indicate that email is permitted as a second factor (Figure 1).
5.6 References to NIST Standards
37% (37) of investigated institutions referenced NIST standards related to authentication security in their online policies (Figure 4). This was most common for Canadian global universities (67%) and US national universities (55%) and least common for US regional colleges (19%).
In some cases, these references acknowledge some noncompliance. For example,
[Anonymized] administration recognizes that fully implementing all controls within the NIST Standards is not possible due to institution limitations and resource constraints. (US regional college)
Others explicitly claim that the institution will adhere to the standard. For example,
[Anonymized] University will: Follow the National Institute of Standards and Technology Framework for Improving Critical Infrastructure Cybersecurity (NIST Cybersecurity Framework). (US national university)
While many others just reference NIST as a guiding resource. For example,
This policy framework consists of eighteen (18) separate policy statements, with supporting Standards documents, based on guidance provided by the National Institute of Standards and Technology. (US liberal arts college)
Figure 4: Whether 101 higher education institutions reference NIST standards related to authentication security in their online policies.
@@ 6 Discussion
The results of this study reveal an alarmingly large gap between many higher education institutions’ authentication policies and the NIST Digital Identity Guidelines. In this section, we discuss the implications of these results.
@@
Our results are particularly worrisome given the increasing prevalence of cyberattacks on educational institutions. The increased use of remote learning and e-learning platforms has increased the threat surface of higher education [5]. According to Alexei and Alexei [2], the threat of “DoS / DDoS attacks, cross-site scripting, spoofing, unauthorized data access and infection with malicious programs, [and] also the theft of personal data has increased dramatically.” Ransomware [36, 37, 1] and phishing [7, 14, 3] have also become more common, with ransomware attack prevalence increasing over 70% in 2023 [12] and recovery costs averaging around \$1 million [35].
The continued use of insecure authentication practices makes institutions even more susceptible to cyberattacks. The NIST compliance rates we observe for higher education institutions is considerably lower than that observed by Hall et al. [24] across a swathe of industries identified as prone to data breaches. This should be a wake up call to higher education IT departments, administration, and all stakeholders about the still limited adoption of the NIST standards, all of which reflect well known best practices. As noted in Section 3, all of these standards are well within the expertise of IT departments to implement, especially given the availability of relevant code libraries and third-party services.
Differences Across Institutional Categories.
We notice that across the NIST standards we investigate, US regional colleges provide less online documentation than institutions in the other categories (more on this below). However, when considering the institutions that do provide documentation, US regional colleges have compliance rates below US liberal arts colleges, which themselves have compliance rates below US national and Canadian global universities. Of the institutions with policies about password expiration, US regional colleges were also the only category with greater than 50% requiring regular password expiration for all affiliates. US regional colleges were also the only category we examined with fewer than 50% requiring MFA overall.
While our results do not provide a causal explanation for this correlation, regional colleges are typically less well-resourced than their top-ranked university or liberal arts counterparts. This could lead to fewer or more overburdened IT staff, less money for third-party MFA solutions, etc. We expect that other categories of traditionally lower-resourced institutions (e.g., community colleges, HSIs, and HBCUs) also have low rates of compliance and encourage additional research for verification.
This finding is important because it provides direct guidance to focus government and industry cybersecurity initiatives. Meaningful gains in cyberattack threat reduction could be made with investment in improving digital authentication security for lower-resourced higher education institutions already less able to afford expensive attack recovery processes. IT conferences, peer workshops, and other fora should be leveraged to help bring all institutions in line with best practices as quickly as possible.
Limited Policy Availability.
The availability of online documentation about authentication practices varied widely across the standards we examined. Very few institutions had posted any information about their use or disuse of password strength meters or email as a second factor for MFA. Approximately 45% of institutions included something about password expiration or knowledge-based authentication in their online documentation. Over 75% of institutions posted about password composition rules or MFA.
The implications of this variety of posted policies are difficult to interpret. For example, consider an institution that does not have an online policy about MFA. It likely means that the institution does not support MFA and therefore does not need to have a MFA policy. However, it is possible that the university does support MFA, but the instructions for setting up and using the service are only available to authenticated affiliates. As an alternative example, consider an institution that does not have an online policy about password composition rules. In this case, one might assume that the institution does not enforce such rules; however, password composition rules are so widespread that it seems unwise make an inference either way. For example, such an institution might have a password symbol/number/case requirement that only appears when already-authenticated users attempt to change their passwords.
@@
Nevertheless, we did observe one consistent trend in the availability of online policies: Across all standards we examined, a greater percentage of US regional colleges had no publicly available policies than institutions in any other category. This is notable, because digital authentication policies ideally should be publicly available online, periodically updated to reflect current practices, and standardized so that disclosures communicate information sufficient to inform relevant stakeholders. Universal compliant disclosure of such policies would serve as a deterrent to cyberattacks and a motivator for other organizations to follow similar practices. It would also serve an educational purpose, teaching affiliates about the authentication practices they should understand and expect in other aspects of their online lives. We hope that this research incentivizes more institutions to bring their digital authentication practices into compliance with known best practices and provide public notification that they have done so.
@@
To simplify this process, we advocate for the standardization of public-facing authentication policy notices, e.g., via a free template. This would make it simple for institutions to 1) see what is necessary to bring their practices into compliance in order to use the template, and 2) quickly create a public-facing document describing these practices with a widely-recognized, user-friendly, and ideally machine readable format. Such a template should allow easy customization for details unique to an institution (e.g., what specific third-party MFA provider is in use) within constraints ensuring that the resulting notice still describes compliant practices. This would be analogous to the standardization of privacy policies and terms and conditions documents, but hopefully with a greater emphasis on readability. The specifics of this standardization would need to be developed in conversation with higher education institutions, but preliminary versions could be proposed and tested in future academic cybersecurity and human-computer interaction research.
@@
As noted in Section 4, online documentation does not always reflect institutional practice. However, only 4 institutions replied to our outreach with meaningful data confirmation or updates. As we discussed, we do not believe that this limits the results of this research. In most cases, online documentation will match practice. IT departments know that community members look to the documentation when they have questions or issues with authentication, so there is an incentive for it to be accurate.
Furthermore, it is useful to evaluate the publicly available policies of institutions whether or not they exactly align with current practice. These policies signal an institution’s values and priorities. If an institution leaves an outdated, noncompliant policy posted online for affiliates and the general public to view, it signals to other institutions that such practices are still acceptable. This can limit the adoption of new best practices and extend existing vulnerabilities.
@@
Nevertheless, the difficulty we faced confirming actual digital authentication practices reflects the difficulty of research in the higher education security and privacy space. Anecdotal conversations with other higher education cybersecurity and privacy researchers suggest that the difficulty of obtaining data is severely limiting the field. Institutions are hesitant to share details of cybersecurity practices with outsiders even if the secrecy of the practices is irrelevant to actual institutional security (c.f. Kerckhoffs’s principle [33]). One IT security director even contacted the authors to apologise that while they would have liked to share more details about their institution’s practices, they had been forbidden to do so by those higher in administration.
@@
We hope that the resistance towards sharing cybersecurity data with academic researchers abates with continued advocacy. The very same institutions that balk at contributing to academic research vetted by institutional review boards readily contract with private education technology platforms having known privacy and security vulnerabilities [34]. Given the breadth of outdated authentication policies we observe in this study, more effort is needed to incentivize institutions to update their practices and better protect their affiliates and operations.
Governance Considerations: Incentives and Knowledge.

### v3

@@ 4.3 Data Collection
This section describes the data collection process followed for all pairs of investigated institutions and standards.
4.3.1 Policy Identification
Institutions are inconsistent as to how they post information about authentication policies online. We therefore used several methods to locate publicly available policies on a wide variety of official institutional websites, including (but not limited to):
@@ 4.3.1 Policy Identification
This process was intentionally open-ended, reflecting the variety of institutional website structures and information posting practices. Relevant policies most frequently appeared on IT support pages (also referred to as IT “guides,” “help desks,” or “knowledge bases”), information security policies, cybersecurity policies, and IT policies.
If we found multiple policies for a particular institution, we selected institution-wide policies over policies for individual academic departments or administrative units. If we found multiple institution-wide policies from different years, we selected the newest policy at the time of data collection. We excluded policies from websites clearly indicated as over 10 years old.
4.3.2 Policy Coding
We read the identified policies, selected the corresponding code from the codebook for each investigated standard, and recorded a link to the policy website to facilitate reproducibility. For the password expiration standard, we also recorded required or recommended password expiration frequencies. For the multi-factor authentication standard, we also recorded what third-party MFA provider was supported or recommended.
4.3.3 Researcher Roles
The policy identification and coding processes described above were performed independently by two non-overlapping sets of researchers:
@@ 4.3.3 Researcher Roles
    The second author initially performed policy identification and coding for all institutions and standards in the study. Their work was then reviewed by a law student research assistant, who visited all recorded links and highlighted any codes they believed were incorrect. The third and fourth authors resolved all disagreements and selected final codes.
@@
    A group of 25 undergraduate computer science students performed policy identification and coding for all institutions and standards as a part of a class on cybersecurity. These students were trained by the first author, and each student collected data and coded policies for 1–6 institutions.
The first author then reviewed and merged all codes produced by (1) and (2) above, resolving disagreements, updating codes for websites that had changed between the initial coding and May 2025, and recording screenshots or PDFs of the institutional policies for reproducibility.
4.4 Reproducibility
We have posted a reproducibility package in a GitHub repository: https://github.com/NoahApthorpe/apthorpe-soups-2025-data. The repository contains a CSV file listing the institutions, institution categories, links to policy websites, all codes, and all additional information collected for the four authentication standards analyzed in this paper. The repository also contains folders with screenshots or PDF documents from the policy websites that were taken in May 2025. Note that the results in this paper reflect these screenshots and PDF documents, although the institutions’ live websites may have changed since these data were collected.
@@ 4.5 Research Ethics
The Colgate University Institutional Review Board determined that this study was exempt. We only report data from publicly available sources posted online by the institutions we study. We do not collect or report any information about individual affiliates at any institution.
All contributing undergraduate computer science students were informed ahead of time that their work may be included in a publication. Their involvement occurred during a normally scheduled lab period and was incorporated into course learning goals. The students received course participation credit for their involvement and are thanked in the acknowledgments.
@@ 4.6 Limitations
First, it is essential to remember that this study examines institutional policies, and that these policies may or may not represent actual, technologically-enforced institutional practices. For example, a policy might say that multi-factor authentication is required for all institution affiliates, but some affiliates might, in practice, be able to disable MFA via their internal profile settings. Similarly, a policy might state that yearly password expiration is required for all affiliates, but the current IT leadership might not actually enforce regular password cycling. Nevertheless, we expect posted policies to generally align with actual practices, especially since posted policies are often primarily intended as guides for affiliates (e.g., on IT “help” websites). The policies also signal the intentions and values of the institution and may influence the actions of affiliates and peer institutions.
On a practical level, determining actual technically-enforced practices would require institutional affiliation or direct input from institutional representatives, which would be challenging. We attempted to address this limitation by emailing the IT departments at 100 of the institutions included in the study with requests to confirm or correct the information we had collected about their policies. However, we only received four meaningful responses, and those effectively corroborated our existing data. We received nine additional responses claiming that no additional data could be provided.
Second, the 135 institutions we studied represent only a fraction of the higher education sector. We encourage follow-up studies examining authentication practices and other forms of cybersecurity compliance across other countries[51] and institutional categories. The institutions in this paper are at the top end of prestige in their respective categories, which must be remembered when interpreting the results. These institutions, in contrast to less prestigious peers, are more likely to be well-resourced, making them more likely to have sufficient staff with expertise and availability to ensure the institution’s authentication (and other cybersecurity) practices follow accepted guidelines. The fact that we find many cases of noncompliance with NIST standards among these prestigious institutions’ online policies suggests that the overall rate of noncompliance is substantially higher across the higher education space.
Third, this study reflects a snapshot in time and only a subset of the authentication standards in NIST SP 800-63. Follow-up longitudinal studies are necessary to see whether the trajectory of institutional policies and practices is trending toward compliance with these and other NIST standards.
@@ 5 Results
Our results for the four standards and 135 institutions we investigated are shown in Table 1. Additional information, including specific affiliates, password expiration frequencies, policy links, and policy screenshots or PDF documents are available in the following GitHub repository: https://github.com/NoahApthorpe/apthorpe-soups-2025-data. This section summarizes these results.
Figure 1: Online publicly available policies of 135 higher education institutions regarding standards from NIST SP 800-63. Policies about multi-factor authentication (in green) align with the standards. Policies about password expiration, password composition requirements, and knowledge-based authentication (in red) are noncompliant.
5.1 Multi-factor Authentication
The use of MFA is widespread in higher education (Figure 1). 70% (95) of all investigated institutions require MFA for all affiliates according to their online policies. 11% (15) require MFA for some affiliates, and 4% (5) at least recommend MFA. The institutions that require MFA at least for some affiliates do so for students, employees, remote users, privileged accounts, or accounts with access to “sensitive” data. Most of the institutions requiring or recommending MFA use a third-party service, such as Duo or Microsoft Authenticator, to provide MFA functionality.
At least 90% of investigated US national universities, US public colleges and universities, Canadian global universities, and US liberal arts colleges require MFA for some or all affiliates. A considerably lower percentage of US HBCUs and US regional colleges require MFA than institutions in other categories (70% and 55% versus $\geq$90%, respectively).
@@ 5.2 Password Expiration
Requiring or recommending password expiration/cycling is still common across higher education (Figure 1). 27% (36) of all investigated institutions require regular password expiration for all affiliates, 10% (14) require regular password expiration for specific affiliates (e.g., faculty and staff, students, affiliates with passwords shorter than a threshold length, or affiliates without MFA enabled), and 13% (18) recommend (but do not require) regular password expiration. A larger percentage of US public colleges and universities (55%) and US HBCUs (50%) require password expiration than institutions in other categories ($\leq$40%).
One year was the most common password expiration frequency we observed, but specific frequencies ranged from expiration every month to expiration frequencies that vary depending on whether an affiliate has enabled MFA or whether the affiliate is an administrator or a regular user (Table 2). Some institutions did not list an exact expiration frequency, rather recommending password expiration “periodically,” “often,” or “frequently.” These policies do not comply with the NIST 800-63 standard that password expiration should only occur in response to a known breach.
@@
1 year (annually, 365 days, etc.) 32
180 days 6
90 days 6
6 months 5
Periodically 5
@@
Often 2
1 year (user), 180 days (admin) 1
1 year (with MFA), 180 days (without MFA) 1
1 year (with MFA), 120 days (without MFA) 1
180 days (user), 90 days (admin) 1
@@
30 days (sensitive data), 90 days (other data) 1
Once or twice a year 1
Frequently 1
Table 2: Required or recommended password expiration frequencies in the online policies of the investigated institutions.
Some of these institutions purposefully disregard the NIST standard. For example, one institution has a Q&A on “Why Are We Implementing A Password Change Initiative When It Is No Longer Considered Best-Practice?” that states
@@
5.3 Password Composition Rules
Despite the NIST SP 600-83 standard that institutions should not impose password composition rules because they “do not significantly improve the security of selected passwords” [62], the presentation and enforcement of password composition rules remains widespread across higher education (Figure 1). 46% (62) of all investigated institutions require that all affiliates meet minimum password composition rules during the password creation process. 10% (14) require that at least some affiliates meet password composition rules, including administrative accounts, accounts with MFA not enabled, or accounts with passwords shorter than a threshold length. Another 13% (18) recommend (but do not require) that affiliates meet password composition rules. Figure 2 shows a representative example of password composition rules from a US regional college.
[…]
@@ 5.4 Knowledge-based Authentication
The use of security questions (e.g., “What is the name of your first pet?”) during the authentication process remains common across higher education (Figure 1). 26% (35) of all investigated institutions require security questions for all affiliates, 4% (6) require security questions for some affiliates, and an additional 11% (15) recommend (but do not require) security questions as an authentication option. All such policies are out of compliance with the NIST SP 800-63 standard that security questions are no longer recommended for any authentication process due to their inherent insecurity. Overall, a larger percentage of US public colleges and universities (55%) and US HBCUs (40%) require security questions than institutions in other categories ($\leq$30%).
@@
The institutions that require security questions typically do not provide any guidance about which questions or what types of answers are likely to be more secure. One institution described transitioning from requiring one security question to requiring three security questions. While this approach may be slightly more secure, it is still out of compliance with the NIST standard.
@@ 6 Discussion
This section discusses the implications of our results, emphasizing insights for the usable security community.
@@
The increased use of remote learning and e-learning platforms has increased the threat surface of higher education [7]. According to Alexei et al. [2], the threat of “DoS / DDoS attacks, cross-site scripting, spoofing, unauthorized data access and infection with malicious programs, [and] also the theft of personal data has increased dramatically.” Ransomware [70, 71, 1] and phishing [12, 24, 3] have also become more common, with ransomware attack prevalence increasing over 70% in 2023 [20] and recovery costs averaging around \$1 million [68].
Given these conditions, the widespread deployment of MFA is encouraging, as the requirement of a second factor substantially decreases the risk of credential-based attacks. However, the persistence of institutions not requiring MFA (especially regional colleges) and the continued use of noncompliant password policies and knowledge-based authentication leaves the sector vulnerable. The NIST SP 800-63 compliance rates we observe for higher education institutions are lower than those observed by Hall et al. [40] across a swath of industries identified as prone to data breaches. Our study should remind higher education IT departments and administrations to require MFA and update outdated password and KBA policies.
Limited Policy Disclosure.
The availability of online documentation about authentication practices varied widely across the institutions we examined. This means that the rates of noncompliance we observe may be a lower bound, with other noncompliant institutions potentially choosing not to post information about their authentication practices online.
@@
Digital authentication policies should be readily available, periodically updated to reflect current practices, and standardized to clearly communicate relevant information. Disclosure of such policies could motivate other organizations to follow similar practices and serve an educational purpose, teaching affiliates about the authentication practices they should understand and expect in other aspects of their online lives. We hope that our research incentivizes more institutions to bring their digital authentication practices into compliance with known best practices and provide public disclosure that they have done so.
@@
To simplify this process, we advocate for the standardization of public-facing authentication policy notices, e.g., via a free template. This would make it simple for institutions to see what changes are necessary to bring their practices into compliance and to quickly create a user-friendly document describing their practices with a widely-recognized and ideally machine readable format. Such a template should allow easy customization for details unique to an institution within constraints ensuring that the resulting notice still describes compliant practices. The specifics of this standardization would need to be developed in conversation with higher education institutions, but preliminary versions could be proposed and tested in future usable security research. This would be similar to prior usable security work on the standardization of privacy policies [44, 21].
@@
The difficulty we faced determining whether actual authentication practices match online policies (Section 4.6) shows that institutions are often unwilling or unable to share details of cybersecurity practices with outside researchers, a phenomenon echoed in the literature [11, 46]. One institution’s IT security director even contacted us to apologize that while they would have liked to share more details about their institution’s practices, they had been forbidden to do so by those higher in administration.
@@
However, the secrecy of details at the level of the NIST standards are irrelevant to the security of an authentication system (see Kerckhoffs’s principle [65] and the ineffectiveness of “security through obscurity”). The very same institutions that balk at contributing to academic research vetted by institutional review boards readily contract with private education technology platforms having known privacy and security vulnerabilities [67]. We hope that the resistance towards sharing cybersecurity data with academic researchers abates with continued advocacy.
Explaining Noncompliance: Incentives.